OpenStudy (anonymous):

Find the least common denominator. Then combine. 2/9 - 2/15=

OpenStudy (amistre64):

spose we set it equal to some arbitrary "n"umber

OpenStudy (anonymous):

can you work it out for me please because i have to show my work and i really dont get it, thankyou

OpenStudy (amistre64):

\[\frac{2}{9}-\frac{2}{15}=N\] now lets just start clearing the fractions from low to high; multiply thru by 9 \[\frac{2(9)}{9}-\frac{2(9)}{15}=(9)N\] \[2-\frac{2(3)}{5}=9N\] now multiply thru by 5 \[2(5)-\frac{2(3)(5)}{5}=9(5)N\] \[2(5)-2(3)=9(5)N\] then solve for N